OpenStudy (anonymous):

a rectangular garden is 10ft long than it is wide. its area is 875ft^2. WHAT ARE ITS DIMENSION? help please >.<

OpenStudy (yrelhan4):

Let width be w.. acc to the question its length, say l will be equal to w+10 Area=w*l=w*(w+10)=875 Solve this equation and find the values of w and l..

OpenStudy (anonymous):

is that w* length or 1 ?

OpenStudy (anonymous):

is it 9.35?

OpenStudy (yrelhan4):

width*length is the area.. your equation would be w^2 + 10w - 875=0.. you just need to solve this quadratic..

OpenStudy (anonymous):

howcome its 10w-875=0?

OpenStudy (anonymous):

shouldn't it equal 875 instead of zero?

OpenStudy (yrelhan4):

w*(w+10)=875 ==> w^2 + 10w=875 ==> w^2 + 10w - 875=0

OpenStudy (anonymous):

and than i do the plus or minus square root of b^2-4ac divided by 2a?

OpenStudy (yrelhan4):

Discriminant D= b^2 -4ac x= [-b (+/-)D]/2a.. this is the quadratic formula..

OpenStudy (anonymous):

yup

OpenStudy (yrelhan4):

you just need to use this and find the answer.